THE ROTATIONAL SPECTRUM OF CHLORINE NITRATE (ClONO_{2}): ν_{6} AND THE ν_{5}/ν_{6}ν_{9} DYAD

\underline{Z. KISIEL}, E.BIALKOWSKA-JAWORSKA, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland; R. A. H. BUTLER, Department of Physics, Pittsburg State University, Pittsburg, KS 66726; D. T. PETKIE, Department of Physics, Wright State University, Dayton, OH 45435; P. HELMINGER, Department of Physics, University of South Alabama, Mobile, Al 36688; F. C. DE LUCIA, Department of Physics, The Ohio State University, Columbus, OH 43210.

Chlorine nitrate is an important stratospheric molecule and analyses of extensive measurements of rotational transitions in the ground state and ν_{9} (120 cm^{−1})^{a}, the 2ν_{9}/ν_{7} dyad (262 cm^{−1})^{b}, and the 3ν_{9}/ν_{7}ν_{9} dyad (361 cm^{−1})^{c} have already been reported. The available experimental data for ClONO_{2} have been extended by a new FASSST spectrum recorded between 118-378 GHz at greater sensitivity and resolution than hitherto and on a more pure sample.

The new spectrum allowed an improvement in the spectroscopic constants for the ground state and ν_{9}, as well as more confident assignment of rotational transitions in higher vibrational states. Application of the AABS package for Assignment and Analysis of Broadband Spectra^{d} was crucial in keeping track of the transitions already assigned and rapid extension of datasets for new states. The analyses of the relatively isolated ν_{6} (435 cm^{−1}) and of the ν_{5}/ν_{6}ν_{9} dyad (551 cm^{−1}) have now been completed for both ^{35}ClONO_{2} and ^{37}ClONO_{2}. In the parent isotopologue the ν_{5} and ν_{6}ν_{9} states are found to be 4.124652(6) cm^{−1} apart and to be connected by a sufficiently strong Coriolis interaction to give rise to measurable interstate transitions and to perturbed hyperfine splitting patterns.

aH.S.P.Müller, et al., J. Mol. Spectrosc., 181, 363-378 (1997).

bR.A.H.Butler et al. J. Mol. Spectrosc., 213, 8-14 (2002).

cR.A.H.Butler et al. J. Mol. Spectrosc., 220, 150-152 (2003).

dZ.Kisiel, PROSPE - Programs for ROtational SPEctroscopy, http://info.ifpan.edu.pl/˜ kisiel/prospe.htm